# PartSDF: Part-Based Implicit Neural Representation for Composite 3D Shape Parametrization and Optimization

### v1

@@ 4.2 Shape Generation
Figure 6: Shape manipulation. We manipulate two shapes per dataset, first by changing the latent of specific parts (car body, mixer helix, and chair backrest) and second by editing part poses (car wheels, mixer width, chair width and height). In all cases, the parts adapts to the modifications and to each other, maintaining a coherent whole with the new parts.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Init. Ours Ours-1P
Figure 7: Part Optimization. (Init.) Initial shape. (Ours) Car body optimized to reduce aerodynamic drag with unchanged wheels. Our part-based representation enables this. (Ours-1P) When using a part-agnostic approach, everything changes, including the wheels. We highlight individual parts in the insert, give the drag coefficient $C_{d}$, and show the surface pressures (from blue to red).
4.3 Part Manipulation
We evaluate our model’s ability to perform part-specific shape manipulation by editing part latent vectors and poses in Figure 6. Both latents and poses can be edited conjointly in PartSDF, we do so separately here for visualization purposes. When modifying part latents, the appearance of the corresponding parts is changed but the shape’s structure and parts layout remain fixed. On the other hand, when changing poses, the parts layout adapts and their general appearance is unchanged. In all cases, the resulting composite shape preserves its overall consistency while fitting the editions. These qualitative results emphasize the flexibility of our part-based representation, showing that our model can adapt individual parts independently while preserving overall shape integrity.
@@ 4.4 Part Optimization
To demonstrate the effectiveness of PartSDF as a part-aware shape prior for downstream tasks, we address a key engineering problem: Refining the shape of a car to reduce the drag induced by air flowing over its surface $\mathcal{S}$, without editing the wheels. Drag can be computed as the surface integral of the air pressure
@@
$\mathcal{L}_{\text{drag}}$ $=C_{d}\left(\mathcal{S}\right)+\mathcal{L}_{\text{reg}}\left(\mathbf{Z},\mathbf{P}\right)\;,$ (10)
$\mathcal{S}$ $=\text{MC}(f_{\theta},\mathbf{Z},\mathbf{P})\;,$
with respect to $\left(\mathbf{Z},\mathbf{P}\right)$, where $\mathcal{L}_{\text{reg}}$ are optional regularization terms and MC is the Marching Cubes algorithm (Lewiner et al., 2003). Importantly, gradients can be computed through the meshing step (Remelli et al., 2020). In practice, we minimize $\mathcal{L}_{\text{drag}}$ over the PartSDF latents $\mathbf{Z}$ and pose parameters $\mathbf{P}$, starting from their values for an existing car. In the example of Fig. 7, we optimize the shapes of several cars while keeping their wheels unchanged, which is something that our part-based approach allows and can not be done with a global representation such as the one of Ours-1P or 3DShape2VecSet (Zhang et al., 2023). While they may achieve lower final drag, the shapes don’t respect design constraints, which is undesirable. For instance, in the second and last row of Figure 7, results with Ours-1P tend to converge to an average car, ignoring the specificity of the initial cars and their wheels. We provide more details and results in Appendix C.
@@ 5 Limitations
The main limitation of PartSDF is its reliance on part labels during training. For our intended application, computer-aided design, this is not a major issue because the decomposition of shapes into parts is known a priori. Thus PartSDF is particularly relevant in scenarios where objects are naturally created with part-aware structures. While this limits PartSDF’s applicability to datasets without predefined decompositions—such as most of those available online—it better matches with practical engineering requirements. Nonetheless, if generalization to unlabeled data is desired, PartSDF could be combined with co-segmentation methods to generate pseudo-labels for its training.
@@ 6 Conclusion
In this work, we introduced PartSDF, a modular and supervised approach specifically designed for composite shape representation. PartSDF enables flexible, part-based shape modeling, supporting independent part manipulation and optimization while maintaining overall shape coherence. Our method leverages a simple architecture, achieving strong performance across tasks such as shape reconstruction, manipulation and generation.
Experimental results demonstrate that PartSDF consistently outperforms baseline methods in part-level accuracy, and can be adapted to a wide range of tasks, highlighting its effectiveness as a robust shape prior for composite structures. This flexibility makes PartSDF particularly suited for applications in fields like engineering design, where precise control over individual components is essential for customization and optimization.
While PartSDF achieves promising results, future work will explore enhancing part interactions to further support applications involving highly dynamic shapes or complex inter-part dependencies. Furthermore, developing more advanced topological supervision for each part should also be investigated to help prevent topologically inconsistent predictions, further improving its the robustness and applicability.
Impact Statement
This paper presents work whose goal is to advance the fields of Machine Learning and Computer-Aided Design. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Table of Contents
In this supplementary material, we provide details regarding our implementation and experimental setup in Appendix B, explain the shape optimization experiment and give additional results in Appendix C, show parts interpolation in Appendix D, ablate components of PartSDF in Appendix E, and we provide additional experimental results in Appendix F: further visualizations for shape reconstruction (Section F.1), generation (Section F.2) and manipulation (Section F.3).
@@ Appendix B Implementation Details
In this section, we describe the architecture of PartSDF (Section B.1), its training procedure (Section B.2), and the point cloud encoder and diffusion models we used (Section B.3). Then, we detail our datasets and their processing (Section B.4), the baselines against which we compare and any hyperparameter choices (Section B.5), and finally the metrics used in this work and how they were computed (Section B.6).
@@ B.1 Network Architecture
The dimensionality of the part latent space is $Z=256$, such that $\mathbf{z}_{p}\in^{256}\;\forall p$, and the hidden size of the fully-connected layers $h_{\text{fc}}$ is 512, with the exception of the model trained on the Chair dataset where it is 256. This is because there are more parts, thus we expect each part to be simpler and it helps to alleviate the memory increase. We use weight normalization (Salimans & Kingma, 2016) on these layers, as in (Park et al., 2019). PartSDF has 8 $h_{\text{fc}}$ layers, counting the input and output ones, with a convolutional layer $h_{\text{conv}}$ between each pair, as described in Section 3.1. All layers, except the output one, are followed by ReLU non-linearities. See Figure 8 for a visualization of their arrangement in our decoder. Note that each $h_{\text{conv}}$ has only $P^{2}+P$ trainable parameters, e.g., $25+5=30$ parameters in the case of cars.
Figure 8: Layer arrangement in PartSDF. (Top) A convolutional layer $h_{\text{conv}}$ is put between each pair of fully-connected layers $h_{\text{fc}}$, with an added residual connection, and every layer is followed by a non-linearity $\sigma$. (Bottom) For a given query position, the input tensor is of shape $P\times D$, with $P$ the number of parts and $D$ the feature size. $h_{\text{fc}}$ operates on the $D$ features for each part separately, while $h_{\text{conv}}$ (with kernels of size $1$) operates on the parts for each feature separately. The output is of shape $P\times D^{\prime}$, with $D^{\prime}$ the output feature size of the second $h_{\text{fc}}$ layer.
@@ B.2 Training Procedure
To supervise the SDF, we follow the sampling strategy used in DeepSDF (Park et al., 2019): 95% of the sampled points are near the surface, and 5% are uniformly distributed in space. The SDF values are computed using the IGL library (Jacobson et al., 2018), assuming watertight meshes. During training, the SDF values are clamped between $-0.1$ and $0.1$ to focus the capacity of the network around the surface.
@@ B.3 Encoder and Diffusion Models
In Section 4.1, we use a point encoder to map point clouds of the full shapes to PartSDF parameter space, i.e. part latents and poses. For this, we leverage the model of 3DShape2VecSet (Zhang et al., 2023) using their official implementation [^1] https://github.com/1zb/3DShape2VecSet. We use a smaller version of the model with 8 self-attention layers (against 24 in the original work), and without the final cross-attention layer with query positions (see Fig. 3 in their work) as we want to output part latents and poses instead of implicit values. We use learnable queries instead of point queries (see Fig. 4(a) in (Zhang et al., 2023)) to encode the point cloud, which creates an order on the latent set, allowing us to recover the parameters of specific parts. Note that we concatenate latent and pose for every part, such that the encoder outputs them both in a single vector.
@@
In order to generate composite shapes with PartSDF, we use a diffusion model to generate the part latents and poses (Section 4.2), more specifically SALAD (Koo et al., 2023) with their official implementation [^2] https://github.com/KAIST-Visual-AI-Group/SALAD. For the second diffusion on the latents, we however make use of a transformer decoder instead of encoder so that each part latent can attend to all part poses, and the part latents are normalized before training. This last point was found crucial for the diffusion as our latent vectors tend to have significantly lower norms than random vectors sampled from a multivariate Gaussian distribution. Indeed, as in (Park et al., 2019), latent vectors are initialized such that $\mathbb{E}_{\mathbf{z}_{p}}\lVert\mathbf{z}_{p}\rVert=1$, while Gaussian noise in $d$-dimensions has $\mathbb{E}_{\mathbf{x}\sim\mathcal{N}^{d}}\lVert\mathbf{x}\rVert=\sqrt{d}$.
@@ B.4 Datasets
For all datasets, we preprocess the meshes by centering them using their bounding box and rescaling them into the cube $[-0.9,0.9]^{3}$, ensuring the longest edge of the bounding box measures $1.8$. This leaves a margin around the shape to prevent the surface from leaving the meshing region, defined as the $[-1,1]^{3}$ cube.
To create the SDF supervision, we perform two types of sampling (Park et al., 2019). First, we sample points on the surface of the mesh and perturb them with Gaussian noise, using variances of $0.005$ and $0.0005$ to generate points near the surface. Second, we uniformly sample points within the $[-1,1]^{3}$ cube. The SDF values for all sampled points are computed using libigl (Jacobson et al., 2018). The samples and their SDF are mixed in a 95/5% ratio (near-surface to uniform samples) and stored for training. During training, a subset of these precomputed points is used, randomized at each epoch.
@@
As explained in Section 4, obtaining high-quality composite shape data is challenging, with online data usually containing non-watertight meshes, or without part decomposition or an inconsistent one. Therefore, we use three curated datasets in this work.
@@
In order to perform Computational Fluid Dynamics (CFD) simulations on the cars, see Appendix C below, we need them to be correctly closed and physically plausible, i.e. wheels and body correctly separated. We therefore hand-process a subset of ShapeNet (Chang et al., 2015)’s cars: The wheels and body are manually selected and separated, and then remeshed to become watertight (Wang et al., 2022).
@@
Each dataset is split into train and test sets ($80\%/20\%$). Model hyperparameters are tuned using the train and validation sets, and final models, as reported below, are trained on train+validation and evaluated on the test set. Part poses are obtained by fitting each part with a cuboid or a cylinder.
@@ B.5 Baselines
Similarly, BAE-Net (Chen et al., 2019) encodes a voxelized input to generate part implicit fields, which can be used for co-segmentation. The key difference, and the reason for comparison to our approach, is that BAE-Net also introduces a weak supervision scheme where ground truth segmentations are provided for a small subset of the training shapes (typically ranging from 1 to 8 shapes). In this setting, the model propagates the learned decomposition across the entire dataset. Similar to DAE-Net, BAE-Net suffers from the limitations of voxelized inputs, which restrict its reconstruction accuracy. However, as shown in Figure 4, the weak supervision allows it to recover the desired decomposition in most cases, although certain parts, such as car wheels, are sometimes segmented together rather than individually. We use the official implementation [^4] https://github.com/czq142857/BAE-NET to train the models. For the weak supervision, we experiment with 1, 2, 3, and 8 shots, finding that 8-shot supervision consistently yields the best results. Therefore, we use these models.
@@
As a fully supervised baseline, we compare against PQ-Net (Wu et al., 2020). PQ-Net employs a two-stage training process: first, an implicit auto-encoder is trained to reconstruct all individual parts, followed by a GRU-based RNN (Cho et al., 2014) that learns to auto-encode sequences of parts, represented by their latent vectors and bounding box parameters. For shape generation, latentGANs (Achlioptas et al., 2018) are used to sample in the RNN’s latent space. Like the previous baselines, PQ-Net relies on voxelized data and often reconstructs ”inflated” shapes. While this helps against thinner parts, such as mixer helices and chair legs, the overall reconstruction accuracy remains low. Additionally, backpropagating through RNNs is notoriously difficult (Pascanu et al., 2013), and the non-continuous nature of its sequence generation makes it less suitable as a shape prior for optimization tasks. We train PQ-Net using the author’s official implementation [^5] https://github.com/ChrisWu1997/PQ-NET.
@@
In order to evaluate the reconstruction and generation quality of our shapes against a more powerful baseline, we also compare to a state-of-the-art part-agnostic method, namely 3DShape2VecSet (Zhang et al., 2023). The core of the architecture is a large auto-encoder transformer (Vaswani et al., 2017) model that encodes a point cloud into an unordered set of latent vectors, which are then decoded at queried positions to obtain occupancy values. For generation, a KL-divergence loss is added during training to make the auto-encoder variational and help a second-stage diffusion model. This diffusion model is trained to generate sets of latent vectors, thus sampling 3DShape2VecSet’s latent space. As noted in the work, the models are very heavy, requiring multiple GPUs for training. Using the author’s official implementation [^6] https://github.com/1zb/3DShape2VecSet, we train auto-encoders without the KL-divergence loss for reconstruction experiments and with the KL-divergence for the generation experiments (as reference, we use the same PartSDF decoder for both experiments). We use the provided hyper-parameters, except for the batch size that we limit to $256$, from $512$, because of memory constraints.

### v2

@@ 4.3 Part Manipulation
Figure 7: Part manipulation. We manipulate two shapes per dataset, first by changing the latent of specific parts (car body, mixer helix, and chair backrest) and second by editing part poses (car wheels, mixer width, chair width and height). In all cases, the parts adapt to the modifications and to each other, maintaining a coherent whole with the new parts.
We evaluate our model’s ability to perform part-specific shape manipulation by editing part latent vectors and poses in Figure˜7. Both latents and poses can be edited conjointly in PartSDF; we do so separately here for visualization purposes. When modifying part latents, the appearance of the corresponding parts is changed, but the shape’s structure and parts layout remain fixed. On the other hand, when changing poses, the part layout adapts, and their general appearance is unchanged. In all cases, the resulting composite shape preserves its overall consistency while fitting the editions. These qualitative results emphasize the flexibility of our part-based representation, showing that our model can adapt individual parts independently while preserving overall shape integrity.
@@ 4.4 Part Optimization
Initial car Ours Ours-1P
Refer to caption Refer to caption0.635 Refer to caption Refer to caption0.413 Refer to caption0.383
Refer to caption Refer to caption0.466 Refer to caption Refer to caption0.361 Refer to caption0.348
Refer to caption Refer to caption0.497 Refer to caption Refer to caption0.473 Refer to caption0.308
Figure 8: Part-based optimization. We refine the car bodies to reduce aerodynamic drag while preserving the size and position of the wheels. (Ours) Our part-based representation enables this. (Ours-1P) In contrast when using a single-part, everything changes, including the wheels. We highlight individual parts on the left, the surface pressure on the right (from blue to red), and give the drag coefficient $C_{d}$ as an inset.
To demonstrate the effectiveness of PartSDF as a part-aware shape prior for downstream tasks, we address a key engineering problem: Refining the shape of a car to reduce the drag induced by air flowing over its surface $\mathcal{S}$, without editing the wheels. Drag can be computed as the surface integral of the air pressure:
@@
$\mathcal{L}_{\text{drag}}$ $=C_{d}\left(\text{MC}(f_{\theta},\mathbf{Z},\mathbf{P})\right)+\mathcal{L}_{\text{reg}}\left(\mathbf{Z},\mathbf{P}\right)\;,$ (11)
with respect to $\left(\mathbf{Z},\mathbf{P}\right)$, where $\mathcal{L}_{\text{reg}}$ are optional regularization terms and MC is the Marching Cubes algorithm (Lewiner et al., 2003). Importantly, gradients can be computed through the meshing step (Remelli et al., 2020). In practice, we minimize $\mathcal{L}_{\text{drag}}$ over the PartSDF latents $\mathbf{Z}$ and pose parameters $\mathbf{P}$, starting from their values for an existing car. In the example of Fig. 8, we optimize the shapes of several cars while keeping their wheels unchanged, which is something that our part-based approach allows and can not be done with a global representation such as the one of Ours-1P or 3DShape2VecSet (Zhang et al., 2023). While they may yield a lower final drag, the shapes do not respect design constraints, which is undesirable. For instance, in the second and last row of Figure˜8, results with Ours-1P tend to converge to an average car, ignoring the specificity of the initial cars and their wheels. We provide more details and results in Appendix˜C.
4.5 Ablation Study
Table 3: Ablation on PartSDF components. We ablate the use of part poses, latent modulation (latMod), and cross-part layers ($h_{\mathrm{cp}}$) on the Car dataset. Best performance is achieved with poses and latent modulation, though we note that $h_{\mathrm{cp}}$ layers decrease slightly the part accuracy of our model. However, this is counterbalanced by enabling inter-part adaptability; see Figure˜9.
CD ($\downarrow$) IoU (%, $\uparrow$) IC ($\uparrow$) pIoU (%, $\uparrow$)
w/o poses 1.43 97.81 0.928 91.71
w/o latMod 1.39 97.37 0.923 92.33
w/o $h_{\mathrm{cp}}$ 1.27 98.02 0.932 96.37
PartSDF 1.27 98.02 0.931 94.89
Refer to caption Refer to caption
w/o $h_{\mathrm{cp}}$ w/ $h_{\mathrm{cp}}$
Refer to caption Refer to caption Refer to caption
GT w/o $\mathcal{L}_{\text{inter}}$ w/ $\mathcal{L}_{\text{inter}}$
(a) Cross-part layers (b) Non-intersection loss
Figure 9: Ablation study of PartSDF. (a) Without the cross-part layers $h_{\mathrm{cp}}$ (Section˜3.2), manipulating a part would not affect the others: The translated and scaled wheels incorrectly collide with the car. With these layers, manipulating some parts correctly affects the others, and the car adapts to the new wheels without intersection. (b) Without the non-intersection loss $\mathcal{L}_{\text{inter}}$, the parts incorrectly overlap. With it, the correct decomposition is learned. We visualize an exploded view of the predicted parts to see the overlaps.
We conduct an ablation study to evaluate the impact of cross-part layers $h_{\mathrm{cp}}$, the use of pose parameters, and latent modulation in PartSDF, and report metrics on the Car dataset in Table˜3. Pose parameters allow independent control of part transformations, which improves the shape reconstruction performance but also allows more precise part manipulation. Latent modulation, i.e., injecting the part latent vectors $\mathbf{z}_{p}$ at every $h_{\mathrm{sp}}$ layer, is a simple modification that improves surface and part reconstruction, as measured by Chamfer-Distance and part IoU. On the other hand, cross-part layers $h_{\mathrm{cp}}$ cause a slight decrease in part accuracy, largely counterbalanced by the inter-part adaptability it enables. In Figure˜9, we show that adding these layers to PartSDF maintains the consistency between parts when doing part manipulation. We also show the necessity of the non-intersection loss $\mathcal{L}_{\text{inter}}$: without it, the losses do not prevent the parts from bleeding into one another, causing the model to learn an incorrect part decomposition with overlapping parts.
@@ 5 Limitations
The main limitation of PartSDF is its reliance on part labels during training. For our intended application, computer-aided design, this is not a major issue because shapes are typically constructed from individual parts; hence, the decomposition of shapes into parts is known a priori. Thus, PartSDF is particularly relevant in scenarios where objects are naturally created with part-aware structures. While this limits PartSDF’s applicability to datasets without predefined decompositions—such as most of those available online—it better matches with practical engineering requirements. Nonetheless, if generalization to unlabeled data is desired, PartSDF could be combined with co-segmentation methods to generate pseudo-labels for its training. Additionally, while we devised a strategy to handle non-watertight parts, it still assumes that they are volumetric and can be represented by an SDF. Thus, parts that are pure surfaces, such as car hoods, would not be naturally handled without inflating them by wrapping a volume around them. An interesting direction for future work would be to use unsigned distance functions to solve this.
@@ 6 Conclusion
In this work, we introduced PartSDF, a modular and supervised approach specifically designed for composite shape representation. PartSDF enables flexible, part-based shape modeling, supporting independent part manipulation and optimization while maintaining overall shape coherence. Our method leverages a simple architecture, achieving strong performance across tasks such as shape reconstruction, manipulation, and generation. Experimental results demonstrate that PartSDF consistently outperforms baseline methods in part-level accuracy and can be adapted to a wide range of tasks, highlighting its effectiveness as a robust shape prior for composite structures. This flexibility makes PartSDF particularly suited for applications in fields like engineering design, where precise control over individual components is essential for customization and optimization. While PartSDF achieves promising results, future work will explore enhancing part interactions to further support applications involving highly dynamic shapes or complex inter-part dependencies. Furthermore, developing more advanced topological supervision for each part should also be investigated to help prevent topologically inconsistent predictions, further improving its robustness and applicability.
@@ Appendix A Table of Contents
In this supplementary material, we provide details regarding our implementation and experimental setup in Appendix˜B, explain the shape optimization experiment and give additional results in Appendix˜C, show parts interpolation in Appendix˜D, and we provide additional experimental results in Appendix˜E: further visualizations for shape reconstruction (Section˜E.1), generation (Section˜E.2) and manipulation (Section˜E.3).
@@ Appendix B Implementation Details
In this section, we describe the architecture of PartSDF (Section˜B.1), its training procedure (Section˜B.2), and the point cloud encoder and diffusion models we used (Section˜B.3). Then, we detail our datasets and their processing (Section˜B.4), the baselines against which we compare and any hyperparameter choices (Section˜B.5), and finally the metrics used in this work and how they were computed (Section˜B.6).
@@ B.1 Network Architecture
The dimensionality of the part latent space is $Z=256$, such that $\mathbf{z}_{p}\in\mathbb{R}^{256}\;\forall p$, and the hidden size of the single-part layers $h_{\mathrm{sp}}$ is 512, with the exception of the model trained on the Chair dataset where it is 256. This is because there are more parts, thus we expect each part to be simpler and it helps to alleviate the memory increase. We use weight normalization (Salimans & Kingma, 2016) on these layers, as in (Park et al., 2019). PartSDF has 8 $h_{\mathrm{sp}}$ layers, counting the input and output ones, with a cross-part layer $h_{\mathrm{cp}}$ between each pair, as described in Section˜3.2. All layers, except the output one, are followed by ReLU non-linearities. Note that each $h_{\mathrm{cp}}$ has only $P^{2}+P$ trainable parameters, e.g., $25+5=30$ parameters in the case of cars.
@@ B.2 Training Procedure
To supervise the SDF, we follow the sampling strategy used in DeepSDF (Park et al., 2019): 95% of the sampled points are near the surface, and 5% are uniformly distributed in space. The SDF values are computed using the IGL library (Jacobson et al., 2018), assuming watertight meshes. During training, the SDF values are clamped between 0.1-0.1 and $0.1$ to focus the capacity of the network around the surface.
@@ B.3 Encoder and Diffusion Models
In Section˜4.1, we use a point encoder to map point clouds of the full shapes to PartSDF parameter space, i.e. part latents and poses. For this, we leverage the model of 3DShape2VecSet (Zhang et al., 2023) using their official implementation [^1] https://github.com/1zb/3DShape2VecSet. We use a smaller version of the model with 8 self-attention layers (against 24 in the original work), and without the final cross-attention layer with query positions (see Fig. 3 in their work) as we want to output part latents and poses instead of implicit values. We use learnable queries instead of point queries (see Fig. 4(a) in (Zhang et al., 2023)) to encode the point cloud, which creates an order on the latent set, allowing us to recover the parameters of specific parts. Note that we concatenate latent and pose for every part, such that the encoder outputs them both in a single vector.
@@
In order to generate composite shapes with PartSDF, we use a diffusion model to generate the part latents and poses (Section˜4.2), more specifically SALAD (Koo et al., 2023) with their official implementation [^2] https://github.com/KAIST-Visual-AI-Group/SALAD. For the second diffusion on the latents, we however make use of a transformer decoder instead of encoder so that each part latent can attend to all part poses, and the part latents are normalized before training. This last point was found crucial for the diffusion as our latent vectors tend to have significantly lower norms than random vectors sampled from a multivariate Gaussian distribution. Indeed, as in (Park et al., 2019), latent vectors are initialized such that $\mathbb{E}_{\mathbf{z}_{p}}\lVert\mathbf{z}_{p}\rVert=1$, while Gaussian noise in $d$-dimensions has $\mathbb{E}_{\mathbf{x}\sim\mathcal{N}^{d}}\lVert\mathbf{x}\rVert=\sqrt{d}$.
@@ B.4 Datasets
For all datasets, we preprocess the meshes by centering them using their bounding box and rescaling them into the cube $0.9,0.9]^{3}$, ensuring the longest edge of the bounding box measures $1.8$. This leaves a margin around the shape to prevent the surface from leaving the meshing region, defined as the $1,1]^{3}$ cube.
To create the SDF supervision, we perform two types of sampling (Park et al., 2019). First, we sample points on the surface of the mesh and perturb them with Gaussian noise, using variances of $0.005$ and $0.0005$ to generate points near the surface. Second, we uniformly sample points within the $1,1]^{3}$ cube. The SDF values for all sampled points are computed using libigl (Jacobson et al., 2018). The samples and their SDF are mixed in a 95/5% ratio (near-surface to uniform samples) and stored for training. During training, a subset of these precomputed points is used, randomized at each epoch.
@@
As explained in Section˜4, obtaining high-quality composite shape data is challenging, with online data usually containing non-watertight meshes, or without part decomposition or an inconsistent one. Therefore, we use three curated datasets in this work.
@@
In order to perform Computational Fluid Dynamics (CFD) simulations on the cars, see Appendix˜C below, we need them to be correctly closed and physically plausible, i.e. wheels and body correctly separated. We therefore hand-process a subset of ShapeNet (Chang et al., 2015)’s cars: The wheels and body are manually selected and separated, and then remeshed to become watertight (Wang et al., 2022).
@@
Each dataset is split into train, validation and test sets ($60\%/20\%/20\%$). Model hyperparameters are tuned using the train and validation sets, and final models, as reported in the main text, are trained on train+validation and evaluated on the test set. Part poses are obtained by fitting each part with a cuboid or a cylinder.
@@ B.5 Baselines
Similarly, BAE-Net (Chen et al., 2019) encodes a voxelized input to generate part implicit fields, which can be used for co-segmentation. The key difference, and the reason for comparison to our approach, is that BAE-Net also introduces a weak supervision scheme where ground truth segmentations are provided for a small subset of the training shapes (typically ranging from 1 to 8 shapes). In this setting, the model propagates the learned decomposition across the entire dataset. Similar to DAE-Net, BAE-Net suffers from the limitations of voxelized inputs, which restrict its reconstruction accuracy. However, as shown in Figure˜5, the weak supervision allows it to recover the desired decomposition in most cases, although certain parts, such as car wheels, are sometimes segmented together rather than individually. We use the official implementation [^4] https://github.com/czq142857/BAE-NET to train the models. For the weak supervision, we experiment with 1, 2, 3, and 8 shots, finding that 8-shot supervision consistently yields the best results. Therefore, we use these models.
@@
As a fully supervised baseline, we compare against PQ-Net (Wu et al., 2020). PQ-Net employs a two-stage training process: first, an implicit auto-encoder is trained to reconstruct all individual parts, followed by a GRU-based RNN (Cho et al., 2014) that learns to auto-encode sequences of parts, represented by their latent vectors and bounding box parameters. For shape generation, latentGANs (Achlioptas et al., 2018) are used to sample in the RNN’s latent space. Like the previous baselines, PQ-Net relies on voxelized data and often reconstructs "inflated" shapes. While this helps against thinner parts, such as mixer helices and chair legs, the overall reconstruction accuracy remains low. Additionally, backpropagating through RNNs is notoriously difficult (Pascanu et al., 2013), and the non-continuous nature of its sequence generation makes it less suitable as a shape prior for optimization tasks. We train PQ-Net using the author’s official implementation [^5] https://github.com/ChrisWu1997/PQ-NET.
PASTA.
Li et al. (2024) proposes another supervised method mostly aimed at generating shape in a part-aware manner. It is made of two parts. First, an autoregressive transformer (Vaswani et al., 2017) is trained to output a sequence of part, each represented through their bounding box (or equivalently, their pose). Secondly, a transformer decoder performs a cross attention between these part bounding boxes and the query points to output an occupancy value for each, using a small final MLP. PASTA’s decoder only rely on the part layout, as described by their respective bounding boxes, and as such entangles the part geometry to them. It is unable to output varying geometry without affecting the part poses. Additionally, PASTA outputs a single occupancy per shape and do not produce individual part. We use the official implementation [^6] https://github.com/Vincent-Li-9701/PASTA to train PASTA on our data.
@@
In order to evaluate the reconstruction and generation quality of our shapes against a more powerful baseline, we also compare to a state-of-the-art part-agnostic method, namely 3DShape2VecSet (Zhang et al., 2023). The core of the architecture is a large auto-encoder transformer (Vaswani et al., 2017) model that encodes a point cloud into an unordered set of latent vectors, which are then decoded at queried positions to obtain occupancy values. For generation, a KL-divergence loss is added during training to make the auto-encoder variational and help a second-stage diffusion model. This diffusion model is trained to generate sets of latent vectors, thus sampling 3DShape2VecSet’s latent space. As noted in the work, the models are very heavy, requiring multiple GPUs for training. Using the author’s official implementation [^7] https://github.com/1zb/3DShape2VecSet, we train auto-encoders without the KL-divergence loss for reconstruction experiments and with the KL-divergence for the generation experiments (as reference, we use the same PartSDF decoder for both experiments). We use the provided hyper-parameters.